OpenStudy (lisa123):

I need a quadratic function that has an x intercept of(4,0) And another function that has an x intercept of(-4,0)

OpenStudy (campbell_st):

so let x = 4 be a zero then x -4 is a factor so you just need another factor... then multiply them together to get the standard form or the quadratic

OpenStudy (campbell_st):

for the 2nd function if x = -4 is an intercept or zero, then x + 4 is a factor so same as above... or you could combine both factors.

OpenStudy (lisa123):

how do I get the other factor @campbell_st

OpenStudy (campbell_st):

well you can use the same value... the quadratic is a perfect square so if you want another factor... let an x intercept by (-1, 0) so what factor would you get from this point..?

OpenStudy (lisa123):

(1,0)? @campbell_st

OpenStudy (campbell_st):

well a zero or intercept is at x = -1 so then x + 1 is a factor you can use this for both questions y = (x-4)(x+1) part 2 y = (x +4)(x + 1)

OpenStudy (lisa123):

oh I see

OpenStudy (lisa123):

I got the answer wrong @campbell_st

OpenStudy (lisa123):

it asks for a parabola that opens upward and one that opens downward

OpenStudy (campbell_st):

well you can have a perfect square \[y = (x -4)^2\] is there any other information in the question that you didn't include..?

OpenStudy (campbell_st):

the example of \[y=(x -4)(x+1) = x^2 -3x - 4\] will open upwards, because the coefficient of x^2 is positive.

OpenStudy (lisa123):

yeah I put that and it says its wrong

OpenStudy (lisa123):

Idk why because the x intercept is 4

OpenStudy (lisa123):

Find two quadratic functions, one that opens upward and one that opens downward, whose graphs have the given x-intercepts. (There are many correct answers.) (−4, 0), (4, 0)

OpenStudy (campbell_st):

ok... so the intercepts are x = -4 so x + 4 is a factor x = 4 so x - 4 is a factor so the quadratic is \[y = a(x -4)(x +4)\] or \[y = a(x^2 - 16)\] for upwards make a a positive number and multiply, 1 is probably easiest for downwards pick a negative number for a so maybe a = -1 and then multiply out... that should help you get the correct answer

OpenStudy (lisa123):

ok let me plug it in

OpenStudy (lisa123):

thank you